FRS Home Page
-
–
S3
–
Pictures
–
Drawings
–
Cable Connections – TPC Electronics – Memo for preparation before beam time
Official Manual of TPC (
PDF )
Recent Log and Memo about TPC (2019) (External link)
Present Configuration (as of 2020/Feb/19)
- In a box (pocket) for S3-vacuum chamber (GTS4EG3T).
- Ti 50um foil for both entrance and exit windows.
- Gap (vertical) = 8 cm (This means -3200 V for drift field voltage)
- Mask scintillator was replaced on 25/April/2019
- Now calibration positions should be x = -25, -10, 0, 10, 25 mm and y = -15, -9, -3 mm
- Maximum allowed HV: -3200 V (drift), +1100 V (anode), -700 V (calibration, PMT).
- For heavy ions, anode and PMT voltages should be much lower than the maximum values.
- Presently, the vertical center of TPC31 is shifted by 1cm downward from the middle of the S3 chamber (memo).
- (probably) installed in the reverserd orientation.
Status during previous beam time
2021/March/28--March/29 (U beam, S483)
- Applied voltages: anode HV = +550V, drift = -3200 V, pmt-grid = -250 V
- Calibration: X is done by grid, Y is by SC31 edge. Y=0 means SC31 center.
- See the attached files for details: (
logbook,
X-fit)
- parameters have been set in /u/profi/NewGo4/NewGo4_2021/online/go4_2021/setup/setup_s483U_01.C
2020/June/05--June/08 (proton beam, WASA000)
- Applied voltages: anode HV = +1150 V, drift = -3200 V
- Not used in main runs. No additional calibration was performed.
2020/May/xx--May/xx (Pb beam, S469, very low evenry)
-
-
2020/Feb/28--Mar/16 (Xe beam, S474, S459, S480)
- Applied voltages: anode HV = +800 V (??? to be checked on logbook), drift HV = -3200V, fiber-PMT = -300V
- Used to get the first impression about the beam poisition at S3. Calibration grid (run???).
-
S3TPC was inserted (by mistake) from XX.XX to YY.YY (run xxx--yyy).
2019/Dec/09--Dec/19 (U beam)
- After changing primary beam from 40Ar to 238U.
- 2019/Dec/16, calibration with U238, 850MeV/u primary beam (file 0905)
- Applied voltages: anode HV = +550 V, drift = -3200V, pmt=-300V (-400 V was too high)
2019/Nov/18--Dec/06 (Ar beam)
- Installed at S3 as TPC31.
- Center of the TPC was not at the beam axis in vertical direction.
- Estimated shift is about 32 mm (see logbook p101--105, 22/Nov).
- Bogdan moved up the end position by 2--3 cm (logbook p130, 26/Nov)
- After that, drift time distribution had no sharp cut. (p138, 28/Nov)
- But we still don't know exact y position of this TPC center.
- All raw signals looked fine, but no calibration was done.
- Gap is 8cm, order: Fiber Sci -> TPC.
- Applied voltages: anode HV = +800 V, drift = -2400V, fiber PMT = -400V.
2019/April/1--3 (107Ag, 290MeV/u, primary beam)
- Installed at S4 as TPC43.
- All signals were checked and ok.
- Signal from the calibration scintillator was strange. Only one line (in y) was observed.
- After the experiment, it turned out that the fiber was broken. See picture below.
- In the beginning, we thought the gap was 6 cm, and applied drift voltage of -2400 V, but in reality it was 8 cm.
- Applied HV: -2400(mistake) and -3200 V (drift), +700 V (anode), -400 or -500 V (calibration PMT)
- Calibration was only partially possible because of the problem in the fiber.
- Gap = 8cm, Fiber Sci. -> TPC.
--
NisanYaprakBenal - 2026-10-26
Updated November 07, 2019
Official Manual of TPC (
PDF
)
Recent Log and Memo about TPC (2019)
(External link)
Messagelog
--
SagarRoy - 2021-06-26